5 points Save A QUESTION 17 many mutations would you expect to see in the human genome (6 x 10 a 60 b.600 6,000 if mismatches could not be repaired? Hint Becouse of toutomeric shit induced mismatches, DNA replication inherently can be no more accurate than 104 m d60,000 e. 00,000
media%2F574%2F574544f6-4538-4b04-8773-40
media%2F84a%2F84a07e55-e896-4c04-bec9-00
Answer #1

Answer 17.

the answer is option e.

Explanation:- As replicative DNA polymerases incorporate noncomplementary nucleotides(mismatched bases) with a frequency of approximately 1:10,000 to 1:100,000 or 1 in 104 to 105 (Arana and Kunkel 2010). Therefore the maximum expected number of mutation would be 6*109/1*104 = 6*105 = 600000.

answer 18.

the answer is option b.

as mentioned in the question 18 that 90% accuracy is there in proofreading process, means 10% of the mutation will move to the next generation or say 10% of total mutations will be in the genome.

therefore 10% of 600000 is 60000.

Answer 19.

the answer is option c.

in case of silent mutation, amino acid sequences of the proteins do not change. therefore protein will function normally.

If there is any mutation occurs in DNA, after the transcription, the triplet codon on mRNA still translates for the same amino acid such mutation called silent mutation.

Answer 20.

the answer is ''false''

insertion of 3 bases is less harmful for protein function because inserting the 3 bases will not change the triplet codons on the mRNA but if we inserted 1 base it will change the entire sequence of all triplet codons on mRNA and protein will be altered entirely.

example. AUG UUU UUC AAC AUU GCA CGA AUC CUG CCG GGA (if 3 bases inserted)

AUG UUG UUU CAA CAU UGC ACG AAU CCU GCC GGG A (if 1 base inserted)